OpenStudy (anonymous):

log 8x - log [1+x^(1/2)] = 2 Solve for x

OpenStudy (anonymous):

\[\log_{10}(8x) - \log_{10}(1+\sqrt{x} ) =2 \]

OpenStudy (anonymous):

so i get \[\log_{10}\frac{ 8x }{ (1+\sqrt{x} ) } = 2\]

OpenStudy (ranga):

\[\Large \log(A) - \log(B) = \log(\frac{ A }{ B })\]If \[\Large \log_{10} y = k\]then\[\Large y = 10^{k}\]

OpenStudy (anonymous):

so \[\frac{ 8x }{(1+\sqrt{x} )} = 100\]

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

then i moved the denominator over to the other side and get\[8x = 100+100\sqrt{x}\]

OpenStudy (ranga):

Yes. Bring them all to the left. Let y = sqrt(x)

OpenStudy (anonymous):

\[8x - 100\sqrt{x} -100\]

OpenStudy (ranga):

= 0.

OpenStudy (anonymous):

\[2x - 25\sqrt{x} - 25 = 0\]

OpenStudy (ranga):

Let y = sqrt(x)

OpenStudy (anonymous):

do you mean substitute y in for sqrt(x) ?

OpenStudy (ranga):

Yes. if y = sqrt(x) then y^2 = x

OpenStudy (anonymous):

\[2y ^{2} - 25y - 25 = 0\]

OpenStudy (ranga):

Solve the quadratic equation for y. Then put back sqrt(x) in place of y and solve for x.

OpenStudy (anonymous):

I'd need the quadratic formula right?

OpenStudy (ranga):

yeah.

OpenStudy (anonymous):

I get x = 90.625 and x = -12.5

OpenStudy (ranga):

What are your y values?

OpenStudy (anonymous):

y = 13.4307 and y = -.9307

OpenStudy (ranga):

That is what I got for y. You square it to get x values.

OpenStudy (anonymous):

okay so x = 180.383 and x = -.866

OpenStudy (ranga):

squaring makes it positive: 180.383 and 0.865 Put those back in the original equation and see if it computes to 2.

OpenStudy (anonymous):

Got it, thank you! The 180.383 works.

OpenStudy (ranga):

Good. You are welcome.